OpenStudy (anonymous):

Find the derivative: 4sec^2 (3pit-3) do I use product rule or chain rule?

OpenStudy (anonymous):

\[4\sec ^{2} (3\pi t - 3)\]

OpenStudy (anonymous):

chain rule ?

zepdrix (zepdrix):

Mhm! Chain rule indeed. Outer-most function is the square.

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{4\frac{d}{dt}\sec^2(3\pi t-3)}=4\cdot2\sec(3\pi t-3)\color{royalblue}{\frac{d}{dt}\sec(3\pi t-3)}\]

OpenStudy (anonymous):

ok..that looks good...hmm

OpenStudy (anonymous):

but i need to take the derv of the inside (3pit -3)

OpenStudy (anonymous):

so its 3pi

zepdrix (zepdrix):

That's off of the very last chain, you need to do this chain first.\[\large\rm =4\cdot2\sec(3\pi t-3)\color{royalblue}{\frac{d}{dt}\sec(3\pi t-3)}\]\[\large\rm =4\cdot2\sec(3\pi t-3)\color{orangered}{\sec(3\pi t-3)\tan(3\pi t-3)}\color{royalblue}{\frac{d}{dt}(3\pi t-3)}\]But yes, you're correct, you'll get an extra 3pi from applying chain rule a second time.

OpenStudy (anonymous):

final answer : \[24 \pi \sec ^{2}(3\pi t-3)\tan(3\pi t-3)\]

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

mm yah looks good! c: